OpenStudy (anonymous):

Solve the equation algebraically. ln (y-1) - ln 2 = x + ln x

OpenStudy (anonymous):

for x or y?

OpenStudy (anonymous):

for x, sorry

OpenStudy (anonymous):

x=ln((y-1)/2x)

OpenStudy (nikvist):

\[\ln{(y-1)-\ln{2}}=x+\ln{x}\] \[\ln{\frac{y-1}{2}}=x+\ln{x}\] \[\frac{y-1}{2}=\exp{(x+\ln{x})}=x\cdot e^x\] \[y=2x\cdot e^x+1\]